I can't seem to understand how to do this also I will like to see if you can verify my answer on the second exersize?
1.Evaluate. | –9 | – | 2 |
2. Multiply. Multiply. (-5/8)(40)= -5/8 x 40/1= 200/8= 25
I found the common denominator for 200 and 8 then I divided and then equal 25


Answer by stanbon(48564)

1.Evaluate. | –9 | – | 2 |
= 9 - 2 = 7
--------------------------
2. Multiply. Multiply. (-5/8)(40)= -5/8 x 40/1= 200/8= 25
I found the common denominator for 200 and 8 then I divided and then equal 25
Don't forget the negative sign.
Ans: -25
=============================
Cheers,
Stan H.